OpenStudy (anonymous):

Find the dot product, a • b. a = 6i + 5j, b = -5i + 4j

OpenStudy (anonymous):

<-30, 20>?

OpenStudy (crashonce):

so is this your question? (6i+5j)(-5i+4j)

OpenStudy (anonymous):

yes

OpenStudy (crashonce):

so expand

OpenStudy (bradely):

(6i+5j).(-5i+4j) (6)(-5)+(5)(4)=-30+20=-10

OpenStudy (crashonce):

-30i^2 +10ji-10ji +9j^2

OpenStudy (crashonce):

=-30i^2+9j^2